OpenStudy (anonymous):

Calculating phase shift of cos function? h(t)= a cos b (x + c) + d a= 0.60 b=200π/1233 c=? d= 1.30

OpenStudy (anonymous):

What I have done so far: h(t)= a cos b (t + c) +d 1.96= 0.60 cos 200π/1233 (0 + c ) + 1.30 1.96= 0.60 cos ( 200π/1233 x 0 ) + ( 200π/1233 x c) +1.30 1.96= 0.60 cos 0 + 200π/1233 c + 1.30 0.66= 0.60 cos 200π/1233 c 1.1= cos 200π/1233 c 〖cos〗^(-1) 1.1= θ θ=200π/1233 〖cos〗^(-1) 1.1= 200π/1233

OpenStudy (anonymous):

But I can't do inverse cos of 1.1, it's not real in calc! So what do I do!

hartnn (hartnn):

how did u get h(t) = 1.96 ?

OpenStudy (anonymous):

Oh I picked a point on the curve h=1.96 t=0

hartnn (hartnn):

1.96= 0.60 cos 200π/1233 (0 + c ) + 1.30 1.96= 0.60 cos( 200π/1233c) + 1.30

hartnn (hartnn):

oh yeah, even i m getting 1.1

hartnn (hartnn):

sure, your h(t) is correct ?

hartnn (hartnn):

oh, waits , is it h(t)= a cos [b (x + c) + d]??

OpenStudy (anonymous):

I'm pretty sure it's just h(t)=a cos b (x + c) +d

hartnn (hartnn):

max value of h(t) must be a+d = 1.9 it can't be 1.96

hartnn (hartnn):

because max value of cos is 1

OpenStudy (anonymous):

a+d?

OpenStudy (anonymous):

Yeah well the thing is, this is a function i have to develop based on real data on tides. So every height of each tide is different so we have to average it.